allie123456:

identify an equation in a point-slope form for the line perpendicular to y=3x+5 that passes through (4,-1) A) y-4=-1/3(x+1) B)y+1=3(x-4) C)y+1= -1/3 (x-4) D)y-1=-3(x+4)

Mercury:

point slope equation:\[y-y_{1}=m(x-x_{1})\] first identify the slope of y = 3x + 5 (major hint: compare to y = mx + b form where m is the slope since we want our line to be perpendicular to y = 3x + 5, take the negative reciprocal of the slope you just got from there, plug m, as well as your (x1,y1) point ((4,-1)) into the point slope equation